OpenStudy (anonymous):

2x^2 +2x =12 solve and check

OpenStudy (zehanz):

You can divide by 2:\[x^2+x=6\]Then subtract 6:\[x^2+x-6=0\]Now use the "sum-product method" or how it's called ;) to write this as a product:\[(x.....)(x.....)=0\]You are looking for two numbers with sum 1 (the number of x-terms) and product -6 (the constant term). They are not difficult to find...

OpenStudy (anonymous):

So would the products be -3, 2

OpenStudy (anonymous):

Thanks for your help. :)

OpenStudy (zehanz):

I'd say 3 and -2, so your equation becomes\[(x+3)(x-2)=0\]Which means solutions are x=-3 and x=2!